# Breather-to-soliton transitions and nonlinear wave interactions for the higher-order modified Gerdjikov–Ivanov equation

### v1

@@ Abstract
In this paper, we systematically investigate the intricate dynamics of the breather-to-soliton transitions and nonlinear wave interactions for the higher-order modified Gerdjikov-Ivanov equation. We discuss the transition conditions of the breather-to-soliton and obtain different types of nonlinear converted waves, including the W-shaped soliton, M-shaped soliton, multi-peak soliton, anti-dark soliton and periodic wave solution. Meanwhile, the interactions among the above nonlinear converted waves are explored by choosing appropriate parameters. Furthermore, we derive the double-pole breather-to-soliton transitions and apply the asymptotic analysis method to analyze the dynamics of the asymptotic solitons for the double-pole anti-dark soliton.
@@
Higher-order modified Gerdjikov–Ivanov equation, Breather-to-soliton transitions, Nonlinear wave interactions, Double-pole converted waves, Asymptotic analysis
@@
[1]organization=School of Mathematical Science, addressline=Beihang University, city=Beijing, postcode=102206, country=China
@@ 1 Introduction
It is well-known that solitons, breathers, and rogue waves represent fundamental nonlinear wave phenomena, each exhibiting distinct dynamic behaviors across diverse physical domains. They have attracted extensive attention due to their emergence in numerous nonlinear models involving various fields such as plasmas physics, nonlinear optics, hydrodynamics and Bose-Einstein condensations[2, 3, 4, 5, 6].
@@
Breathers, which describe the growth of the perturbations in finite continuous backgrounds, contain two different types, namely Kuznetsov–Ma breathers and Akhmediev breathers [7, 8]. Recently, the state transition mechanisms between the breathers (or rogue waves) and solitons have been explored in various nonlinear models. Abundant converted waves including multi-peak solitons, periodic waves and anti-dark solitons have been presented starting from the breathers or rogue waves by an appropriate transitions condition. The first case involves the nonlinear models featuring high-order terms that contribute to the state transitions. The mixed nonlinear Schrödinger equation (NLS)[9], fifth-order NLS[10], higher-order generalized NLS[11] and six-order NLS equation[12] have been successively discovered to possess various converted waves and nonlinear wave interactions. The Hirota equation[13, 14, 15], as an extended NLS-type equation with high-order dispersion terms, also exhibits the breather-to-soliton transition phenomenon. Meanwhile, Sasa–Satsuma equation[16] and the high-order Chen–Lee–Liu equation have also been considered[17]. Another case occurs in coupled equations describing multi-field interactions which facilitate the state transitions. The transitions of the breather solution have been progressively discussed in the coupled Hirota equation[18, 19], multi-component AB system[20, 21, 22], coupled Fokas–Lenells system[23] and coupled extended modified Korteweg-de Vries system[24].
In this paper, we focus on the the higher-order modified Gerdjikov–Ivanov equation (HMGI),
$iq_{t}$ $+2iq_{xxx}+\left[(16\delta+4)|q|^{2}-4\delta\right]q_{xx}+(4\delta+1)q^{2}q_{xx}^{*}+(12\delta+3)q^{*2}q_{x}^{2}$
$+\left[(20\delta+2)qq_{x}^{*}-\left(48\delta^{2}+30\delta+\frac{15}{4}\right)i|q|^{4}\right.\left.+\left(24\delta^{2}+6\delta\right)i|q|^{2}-2i\delta^{2}\right]q_{x}$
$-\left[(24\delta^{2}+6\delta)|q|^{2}q^{2}-8\delta^{2}q^{2}\right]iq_{x}^{*}-\left(16\delta^{3}+24\delta^{2}+\frac{15}{2}\delta+\frac{5}{8}\right)|q|^{6}q$
$+\left(16\delta^{3}+12\delta^{2}+\frac{3}{2}\delta\right)|q|^{4}q-\left(4\delta^{3}+\delta^{2}\right)|q|^{2}q+2\delta^{3}q=0.$ (1)
@@
where $\delta$ is an arbitrary real number which imposes a significant influence on the self-steepening effect in transmission which is the self-frequency shift of Raman scattering. Guan, Yang and Meng et.al. obtained the high-order rogue wave solutions and their dynamic properties for the HMGI equation by generalized Darboux transformation method[25]. The soliton, breather and new rogue wave solutions of the HMGI euation are reached by Zhao, Wang and Yang through Darboux transformation method[26]. The multi-positon solutions, interaction solutions and the modulation instability for the HMGI equation are exihited by Zhao, Wang and Zhang[27].
@@ 2 Breather-to-soliton transitions
$Q_{1}$ $=\frac{1}{2}iq^{*}q_{xx}+\frac{1}{2}iqq_{xx}^{*}+\frac{1}{16}\left[-8iq_{x}^{*}+(488+12)|q|^{2}q-\left(8\lambda^{2}+16\delta\right)q^{*}\right]q_{x}$
$-3\left[q^{*}\left(\delta+\frac{1}{4}\right)q-\frac{\lambda^{2}}{6}-\frac{\delta}{3}\right]qq_{x}^{*}-6i|q|^{6}\left(\delta^{2}+\frac{\delta}{2}+\frac{5}{96}\right)$
$+4i\left(\delta+\frac{3}{16}\right)\left(\frac{\lambda^{2}}{2}+\delta\right)|q|^{4}-\frac{i\left(\lambda^{2}+\delta\right)^{2}|q|^{2}}{2}+i\lambda^{6}+2i\lambda^{4}\delta+i\lambda^{2}\delta^{2}+i\delta^{3},$
$Q_{2}$ $=2\lambda\bigg{[}-\frac{q_{xx}}{2}+\bigg{[}3i\bigg{(}\delta+\frac{1}{4}\bigg{)}|q|^{2}-\frac{1}{2}i\lambda^{2}-i\delta\bigg{]}q_{x}+\bigg{[}i\delta qq_{x}^{*}+\frac{(\lambda^{2}+\delta)^{2}}{2}$
@@
$Q_{3}$ $=2\lambda\bigg{[}\frac{q_{xx}^{*}}{2}+\bigg{[}3i\bigg{(}\delta+\frac{1}{4}\bigg{)}|q|^{2}-\frac{1}{2}i\lambda^{2}-i\delta\bigg{]}q_{x}^{*}+\bigg{[}i\delta^{*}qq_{x}-\frac{(\lambda^{2}+\delta)^{2}}{2}$
@@
where $\lambda$ is the complex spectral parameter and the zero curvature equation $P_{t}-Q_{x}+PQ-QP=0$ generate original Eq.(1).
We introduce the zeed solution $q[0]=ce^{\mathrm{i}(ax+bt)}$ to Eq.(2) with
$b=$ $-16\delta^{3}c^{6}-24\delta^{2}c^{6}-\frac{15}{2}c^{6}\delta-\frac{5}{8}c^{6}+16\delta^{3}c^{4}+24\delta^{2}c^{4}a+12\delta^{2}c^{4}+24c^{4}\delta a$
$+\frac{3}{2}c^{4}\delta+\frac{15}{4}c^{4}a-4\delta^{3}c^{2}-16\delta^{2}c^{2}a-\delta^{2}c^{2}-12c^{2}a^{2}\delta-6c^{2}\delta a-6c^{2}a^{2}$
@@
and $\Omega_{1}$ is given by replacing (n+1)-th column in $\Omega_{2}$ with
@@
$\psi_{j}=[c_{1}(c\lambda_{j}\mathrm{e}^{B_{j}}+(H_{j}-D_{j})\mathrm{e}^{-B_{j}})+c_{2}(-c\lambda_{j}\mathrm{e}^{-B_{j}}+(H_{j}+D_{j})\mathrm{e}^{B_{j}})]\mathrm{e}^{A},$
$\phi_{j}=[c_{1}((H_{j}-D_{j})\mathrm{e}^{B_{j}}+c\lambda_{j}\mathrm{e}^{-B_{j}})+c_{2}((H_{j}+D_{j})\mathrm{e}^{-B_{j}}-c\lambda_{j}\mathrm{e}^{B_{j}})]\mathrm{e}^{-A},$
$A=\mathrm{i}(ax+bt),~{}~{}D_{j}=2\mathrm{i}\delta c^{2}+\lambda_{j}^{2}\mathrm{i}-a\mathrm{i},$
$H_{j}=\sqrt{-4c^{4}\delta^{2}-4c^{2}\delta\lambda_{j}^{2}+4ac^{2}\delta-c^{2}\lambda_{j}^{2}-\lambda_{j}^{4}+2a\lambda_{j}^{2}-a^{2}},$
$B=\frac{H_{j}x}{2}+H_{j}t\Bigg{\{}\left(4\delta^{2}+3\delta+\frac{3}{8}\right)c^{4}+a^{2}+\left(\lambda_{j}^{2}+2\delta\right)a+\left(\lambda_{j}^{2}+\delta\right)^{2}$
$\quad+\left[\left(-4\delta-\frac{3}{2}\right)a-4\left(\frac{\lambda_{j}^{2}}{2}+\delta\right)\left(\delta+\frac{1}{4}\right)\right]c^{2}\Bigg{\}},$
where $c_{1},c_{2}$ are the real coefficients.
@@
Subsequently, we firstly consider the first-order solution, namely $n=1$. Assume $H_{1}=H_{1,R}+\mathrm{i}H_{1,I}$ and $\lambda_{1}=\alpha_{1}+\mathrm{i}\beta_{1}$, we can find that the first-order solution comprises the hyperbolic function parts $\cosh(V_{1,H}t+H_{1,R}x),~{}\sinh(V_{1,H}t+H_{1,R}x)$ and the trigonometric function parts $\cos(V_{1,T}t+H_{1,I}x),~{}\sin(V_{1,T}t+H_{1,I}x)$, where
$V_{1,H}$ $=W_{1}H_{1,R}-W_{2}H_{1,I},~{}V_{T}=W_{1}H_{1,I}+W_{2}H_{1,R},$
@@
$W_{1}$ $=\frac{1}{4}\Bigg{[}32\left(c^{2}-\frac{1}{2}\right)^{2}\delta^{2}+\Bigl{(}24c^{4}-16\alpha_{1}^{2}c^{2}+16\beta_{1}^{2}c^{2}-32ac^{2}-8c^{2}-16\beta_{1}^{2}$
$+16\alpha_{1}^{2}+16a\Bigr{)}\delta+3c^{4}+\bigl{(}-4\alpha_{1}^{2}+4\beta_{1}^{2}-12a\bigr{)}c^{2}+8\alpha_{1}^{4}+\bigl{(}-48\beta_{1}^{2}+8a\bigr{)}\alpha_{1}^{2}$
$+8\beta_{1}^{4}-8a\beta_{1}^{2}+8a^{2}\Bigg{]},$
$W_{2}$ $=4\alpha_{1}\beta_{1}\Big{(}-2c^{2}\delta+2\delta-\frac{c^{2}}{2}-2\beta_{1}^{2}+2\alpha_{1}^{2}+a\Big{)}.$
where the symbols $R$ and $I$ denote the real part and the imaginary part.
@@
The above two function parts respectively describe the localization and periodicity of the wave’s transverse distribution. In other words, the first-order solution is composed of the soliton parts and the periodic parts. It’s obvious that the velocities of the two parts are $\frac{V_{1,H}}{H_{1,R}}$ and $\frac{V_{1,T}}{H_{1,I}}$, respectively. We can derive the first-order breather solution from Eq.(3) when $\frac{V_{1,H}}{H_{1,R}}\neq\frac{V_{1,T}}{H_{1,I}}$. When $\frac{V_{1,H}}{H_{1,R}}=\frac{V_{1,T}}{H_{1,I}}$, the first-order solution will be converted into a continuous soliton. The breather-to-soliton state transition will be discussed.
2.1 $\frac{V_{1,H}}{H_{1,R}}\neq\frac{V_{1,T}}{H_{1,I}}$

### v2

# Breather-to-soliton transitions and nonlinear wave interactions for the higher-order generalized Gerdjikov–Ivanov equation

@@ Abstract
In this paper, we systematically investigate the intricate dynamics of the breather-to-soliton transitions and nonlinear wave interactions for the higher-order generalized Gerdjikov-Ivanov equation. The transition conditions of the breather-to-soliton are established and the novel nonlinear converted waves, including the W-shaped soliton, M-shaped soliton, multi-peak soliton, anti-dark soliton and periodic wave solution are discussed. Meanwhile, the interactions among the above nonlinear converted waves are explored by choosing appropriate parameters. Furthermore, we derive the double-pole solutions exhibiting breather-to-soliton transitions and employ the asymptotic analysis method to analyze the dynamics of the asymptotic solitons for the double-pole anti-dark soliton. This work deepens the fundamental understanding of nonlinear wave metamorphosis induced by higher-order terms in integrable systems.
@@
Higher-order generalized Gerdjikov–Ivanov equation, Breather-to-soliton transitions, Nonlinear wave interactions, Double-pole converted waves, Asymptotic analysis
journal: XXX
@@
[…]
@@ 1 Introduction
It is well-known that solitons, breathers, and rogue waves represent fundamental nonlinear wave phenomena, each exhibiting distinct dynamic behaviors across diverse physical domains. They have attracted extensive attention due to their emergence in numerous nonlinear models involving various fields such as plasma physics, nonlinear optics, hydrodynamics and Bose-Einstein condensations [2, 3, 4, 5, 6].
Breathers, which describe the growth of the perturbations in finite continuous backgrounds, contain two different types, namely Kuznetsov–Ma breathers and Akhmediev breathers [7, 8]. Recently, the state transition mechanisms between the breathers (or rogue waves) and solitons have been explored in various nonlinear models. The conclusion that breathers and rogue waves can convert into solitons via the Raman effect finds its justification and factual basis in [9, 10].
@@
A wealth of converted wave forms—including multi-peak solitons, periodic waves and anti-dark solitons—have been derived from breathers and rogue waves under appropriate transition conditions. The first case involves the nonlinear models featuring higher-order terms that contribute to the state transitions [11]. The mixed nonlinear Schrödinger equation (NLS) [12], fifth-order NLS [13], higher-order generalized NLS [14] and six-order NLS equation [15] have been successively discovered to possess various converted waves and nonlinear wave interactions. The Hirota equation [16, 17, 18], as an extended NLS-type equation with higher-order dispersion terms, also exhibits the breather-to-soliton transition phenomenon. Meanwhile, Sasa–Satsuma equation [19] and the high-order Chen–Lee–Liu equation have also been considered [20]. Another case occurs in coupled equations describing multi-field interactions which facilitate the state transitions. The transitions of the breather solution have been progressively discussed in the coupled NLS equation [21], coupled Hirota equation [22, 23], multi-component AB system [24, 25, 26], coupled Fokas–Lenells system [27] and coupled extended modified Korteweg-de Vries system [28].
In addition, multi-pole solitons, also known as degenerate solitons, have attracted extensive attention in the academic research [29, 30, 31]. They can describe the interactions among multiple chirped pulses with the same amplitude in the optical fibers and in the quantum physical field. Meanwhile, the super-transparent potential has a strong correlation with multi-pole solitons. Hence, the double-pole breather-to-soliton represents a significant area of research.
The third-type derivative nonlinear Schrödinger equation, which is also called Gerdjikov–Ivanov (GI) equation,
$\mathrm{i}q_{t}+q_{xx}+\frac{1}{2}q^{3}q^{*2}-\mathrm{i}q^{2}q^{*}_{x}=0,$ (1)
describes the evolution of complex-valued wave envelopes, particularly in contexts involving nonlinearity and dispersion coupled with specific gain or loss mechanisms. $q(x,t)$ denotes the complex envelope amplitude. GI equation has been extensively studied in recent decades [32, 33, 34, 35]. When increasing the intensity of incident light to generate shorter pulses, relevant experiments have shown that higher-order dispersion terms must be considered. Therefore, we focus on the the higher-order generalized Gerdjikov–Ivanov equation (HMGI),
$\mathrm{i}q_{t}$ $+2\mathrm{i}q_{xxx}+\left[(16\delta+4)|q|^{2}-4\delta\right]q_{xx}+(4\delta+1)q^{2}q_{xx}^{*}+(12\delta+3)q^{*2}q_{x}^{2}$
$+\left[(20\delta+2)qq_{x}^{*}-\left(48\delta^{2}+30\delta+\frac{15}{4}\right)\mathrm{i}|q|^{4}+\left(24\delta^{2}+6\delta\right)\mathrm{i}|q|^{2}-2\mathrm{i}\delta^{2}\right]q_{x}$
$-\left[(24\delta^{2}+6\delta)|q|^{2}q^{2}-8\delta^{2}q^{2}\right]\mathrm{i}q_{x}^{*}-\left(16\delta^{3}+24\delta^{2}+\frac{15}{2}\delta+\frac{5}{8}\right)|q|^{6}q$
$+\left(16\delta^{3}+12\delta^{2}+\frac{3}{2}\delta\right)|q|^{4}q-\left(4\delta^{3}+\delta^{2}\right)|q|^{2}q+2\delta^{3}q=0,$ (2)
@@
here, $\delta$ is an arbitrary real number that exerts a significant influence on the self-steepening effect during transmission, which corresponds to the self-frequency shift in Raman scattering. Guan, Yang, Meng and Liu obtained the higher-order rogue wave solutions and their dynamic properties for the HMGI equation by generalized Darboux transformation method [36]. The soliton, breather and new rogue wave solutions of the HMGI euation were reached by Zhao, Wang and Yang through Darboux transformation method [37]. The multi-positon solutions, interaction solutions and the modulation instability for the HMGI equation were exhibited by Zhao, Wang and Zhang [38].
@@ 2 Breather-to-soliton transitions
$Q_{1}$ $=\frac{1}{2}\mathrm{i}q^{*}q_{xx}+\frac{1}{2}\mathrm{i}qq_{xx}^{*}+\frac{1}{16}\left[-8\mathrm{i}q_{x}^{*}+(488+12)|q|^{2}q-\left(8\lambda^{2}+16\delta\right)q^{*}\right]q_{x}$
$\quad-3\left[q^{*}\left(\delta+\frac{1}{4}\right)q-\frac{\lambda^{2}}{6}-\frac{\delta}{3}\right]qq_{x}^{*}-6\mathrm{i}|q|^{6}\left(\delta^{2}+\frac{\delta}{2}+\frac{5}{96}\right)$
$\quad+4\mathrm{i}\left(\delta+\frac{3}{16}\right)\left(\frac{\lambda^{2}}{2}+\delta\right)|q|^{4}-\frac{\mathrm{i}\left(\lambda^{2}+\delta\right)^{2}|q|^{2}}{2}+\mathrm{i}\lambda^{6}+2\mathrm{i}\lambda^{4}\delta+\mathrm{i}\lambda^{2}\delta^{2}+\mathrm{i}\delta^{3},$
$Q_{2}$ $=2\lambda\bigg{[}-\frac{q_{xx}}{2}+\bigg{[}3\mathrm{i}\bigg{(}\delta+\frac{1}{4}\bigg{)}|q|^{2}-\frac{1}{2}\mathrm{i}\lambda^{2}-\mathrm{i}\delta\bigg{]}q_{x}+\bigg{[}\mathrm{i}\delta qq_{x}^{*}+\frac{(\lambda^{2}+\delta)^{2}}{2}$
@@
$Q_{3}$ $=2\lambda\bigg{[}\frac{q_{xx}^{*}}{2}+\bigg{[}3\mathrm{i}\bigg{(}\delta+\frac{1}{4}\bigg{)}|q|^{2}-\frac{1}{2}\mathrm{i}\lambda^{2}-\mathrm{i}\delta\bigg{]}q_{x}^{*}+\bigg{[}\mathrm{i}\delta^{*}qq_{x}-\frac{(\lambda^{2}+\delta)^{2}}{2}$
@@
where $\lambda$ is the complex spectral parameter and the zero curvature equation $P_{t}-Q_{x}+PQ-QP=0$ generates original Eq.(1).
We introduce the seed solution $q[0]=de^{\mathrm{i}(ax+bt)}$ to Eq.(3) with
$b=$ $-16\delta^{3}d^{6}-24\delta^{2}d^{6}-\frac{15}{2}d^{6}\delta-\frac{5}{8}d^{6}+16\delta^{3}d^{4}+24\delta^{2}d^{4}a+12\delta^{2}d^{4}+24d^{4}\delta a$
$+\frac{3}{2}d^{4}\delta+\frac{15}{4}d^{4}a-4\delta^{3}d^{2}-16\delta^{2}d^{2}a-\delta^{2}d^{2}-12d^{2}a^{2}\delta-6d^{2}\delta a-6d^{2}a^{2}$
@@
and $\Omega_{1}$ is given by replacing the ($n+1$)-th column in $\Omega_{2}$ with
@@
$\psi_{j}=[c_{1}(d\lambda_{j}\mathrm{e}^{B_{j}}+(H_{j}-D_{j})\mathrm{e}^{-B_{j}})+c_{2}(-d\lambda_{j}\mathrm{e}^{-B_{j}}+(H_{j}+D_{j})\mathrm{e}^{B_{j}})]\mathrm{e}^{A},$
$\phi_{j}=[c_{1}((H_{j}-D_{j})\mathrm{e}^{B_{j}}+d\lambda_{j}\mathrm{e}^{-B_{j}})+c_{2}((H_{j}+D_{j})\mathrm{e}^{-B_{j}}-d\lambda_{j}\mathrm{e}^{B_{j}})]\mathrm{e}^{-A},$
$A=\mathrm{i}(ax+bt),~~D_{j}=2\mathrm{i}\delta d^{2}+\mathrm{i}\lambda_{j}^{2}-\mathrm{i}a,$
$H_{j}=\sqrt{-4d^{4}\delta^{2}-4d^{2}\delta\lambda_{j}^{2}+4ad^{2}\delta-d^{2}\lambda_{j}^{2}-\lambda_{j}^{4}+2a\lambda_{j}^{2}-a^{2}},$
$B=\frac{H_{j}x}{2}+H_{j}t\Bigg{\{}\left(4\delta^{2}+3\delta+\frac{3}{8}\right)d^{4}+a^{2}+\left(\lambda_{j}^{2}+2\delta\right)a+\left(\lambda_{j}^{2}+\delta\right)^{2}$
$\quad\quad+\left[\left(-4\delta-\frac{3}{2}\right)a-4\left(\frac{\lambda_{j}^{2}}{2}+\delta\right)\left(\delta+\frac{1}{4}\right)\right]d^{2}\Bigg{\}},$
$c_{1},c_{2}$ are the real coefficients.
@@
Subsequently, we firstly consider the first-order solution, namely $n=1$. Assuming $H_{1}=H_{1,R}+\mathrm{i}H_{1,I}$, which corresponds to the spectral parameter $\lambda_{1}=\alpha_{1}+\mathrm{i}\beta_{1}$, we can find that the first-order solution comprises the hyperbolic function parts $\cosh(tV_{1,H}+xH_{1,R})$, $\sinh(tV_{1,H}+xH_{1,R})$ and the trigonometric function parts $\cos(tV_{1,T}+xH_{1,I})$, $\sin(tV_{1,T}+xH_{1,I})$, where
$V_{1,H}$ $=W_{1}H_{1,R}-W_{2}H_{1,I},~~~V_{1,T}=W_{1}H_{1,I}+W_{2}H_{1,R},$
@@
$W_{1}$ $=\frac{1}{4}\Bigg{[}32\left(d^{2}-\frac{1}{2}\right)^{2}\delta^{2}+\Bigl{(}24d^{4}-16\alpha_{1}^{2}d^{2}+16\beta_{1}^{2}d^{2}-32ad^{2}-8d^{2}-16\beta_{1}^{2}$
$\quad+16\alpha_{1}^{2}+16a\Bigr{)}\delta+3d^{4}+\bigl{(}-4\alpha_{1}^{2}+4\beta_{1}^{2}-12a\bigr{)}d^{2}+8\alpha_{1}^{4}+\bigl{(}-48\beta_{1}^{2}+8a\bigr{)}\alpha_{1}^{2}$
$\quad+8\beta_{1}^{4}-8a\beta_{1}^{2}+8a^{2}\Bigg{]},$
$W_{2}$ $=4\alpha_{1}\beta_{1}\Big{(}-2d^{2}\delta+2\delta-\frac{d^{2}}{2}-2\beta_{1}^{2}+2\alpha_{1}^{2}+a\Big{)},$
and the symbols $R$ and $I$ denote the real part and the imaginary part.
@@
The above two function parts respectively describe the localization and periodicity of the wave’s transverse distribution. In other words, the first-order solution is composed of the soliton parts and the periodic parts. It’s obvious that the velocities of the two parts are $\frac{V_{1,H}}{H_{1,R}}$ and $\frac{V_{1,T}}{H_{1,I}}$, respectively. We can derive the first-order breather solution from Eq.(4) when $\frac{V_{1,H}}{H_{1,R}}\neq\frac{V_{1,T}}{H_{1,I}}$ holds. If $\frac{V_{1,H}}{H_{1,R}}=\frac{V_{1,T}}{H_{1,I}}$, the first-order solution will be converted into a continuous soliton. The breather-to-soliton state transition will be discussed.
2.1 The case $\frac{V_{1,H}}{H_{1,R}}\neq\frac{V_{1,T}}{H_{1,I}}$
In this case, the first-order solution becomes the localized wave solution with breathing behavior. Under the condition
$\alpha_{1}^{2}-\beta_{1}^{2}+\frac{(4\delta+1)d^{2}-2a}{2}=0,$ (5)
when adding the following condition
$4\beta_{1}^{2}\alpha_{1}^{2}+\frac{d^{4}}{4}+d^{2}\left(2d^{2}\delta-a\right)>0,$ (6)